OpenStudy (anonymous):

The times will never be the same A marble is drawn at random first from a jar containing four black and four white marbles, and then from a jar containing six black and two white marbles. What is the probability of drawing two black marbles? A. 3/8 B. 1/2 C. 5/8

OpenStudy (anonymous):

I think its A

OpenStudy (anonymous):

You sure?

OpenStudy (anonymous):

Why do you think that?

OpenStudy (anonymous):

3/8 is correct

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

P in first case = 1/2 P in second case = 3/4 so final prob will be 1/2 multiplied by 3/4 = 3/8

OpenStudy (anonymous):

will the probability of drawing a black marble from the first jar is 1/2 and the probablility of getting a black from the second jar is 6/8 and when you multiply 1/2 by 6/8 you get 3/8

OpenStudy (anonymous):

(4/8)x(6/8)=3/8

OpenStudy (anonymous):

Which of the following represents the number of subsets of five elements can be formed from a set of six elements? A. 5! B. C(6, 5) C. P(6, 5)